OpenStudy (anonymous):

Solve the system of linear equations below. x + y = 4 2x + 3y = 0 A.x = -6, y = 2 B.x = -1, y = 5 C.x = 11/5, y = 9/5 D.x = 12, y = -8

OpenStudy (anonymous):

x + y = 4 ------- 1) 2x + 3y = 0 ------ 2) multiply 1) by 3 to eliminate y 3x + 3y = 12 ----- 3) subtract 2) from 3) to get x: x = 12 Recall: x + y = 4 y = 4 - 12 = -8 x = 12, y = -8

OpenStudy (anonymous):

thank you